Louisiana State University Health Sciences Center Administration & Finance
 

From: Ronnie Smith, Vice Chancellor, Administration and Finance
Subject: Hurricane Katrina Recovery – Update #32
Posted: December 22, 2006

The architects recently met with representatives of the Dental School and LSUHSC administration to determine what final preparations are necessary to re-occupy the upper floors of the buildings and to return the School to operations in New Orleans for the next academic calendar year. The final environmental remediation, provision of sterilization capabilities, temporary space for Facilities personnel, and Receiving were among the items identified. LSUHSC has taken responsibility for executing the environmental remediation contract and plans are to have work beginning in early February, 2007. The architects are pursing options for providing for the sterilization in ways that FEMA will reimburse. Repairs to air handling units, thermostats and related items throughout the two main buildings continue. Pricing has also been received for a repair to the computerized HVAC monitoring system that will give the limited Facilities staff the tools necessary to monitor basic building conditions and to respond when needed.

The project to improve the lighting at the Student Parking lots on the downtown campus has been completed. All new light poles and fixtures have been installed to provide much improved illumination and a safer environment for the students to park in.

At 1542 Tulane Ave., a variety of repair work continues to be performed. The elevator contractor has increased his presence and has been working some weekends to expedite repairs to both the freight elevator and to the bank of four passenger elevators located at the center rear of the building. The repaired building systems are being tested and leaks are being repaired as they are discovered. Much of the flood damaged switch gear and transformers have been removed from the basement. Old gas pipe has been removed and new sump pumps are being installed in the basement. An additional transformer is being installed and the feed repaired so that the final portion of the building will have electrical power restored.

Here are highlights of work that has been accomplished since our last update was posted two weeks ago.

Project to Make Temporary Repairs to 1542 Tulane Ave.:

  • New conduit for the wiring to connect an emergency generator has been installed.
  • The flood damaged switch gear is being removed from the basement.
    • A specialty contractor removed the oil from the old transformers that were located in the vault under the Lakeside lot in preparation for their removal.
    • The Lakeside lot has been closed for the week of 12/18 through 12/22 to allow a large crane space to operate safely as it removes the switch gear, transformers, and other heavy items from the basement and the Lakeside vault. A new switch gear will also be installed in the basement.
    • The six old transformers are now in the Lakeside lot waiting to be picked up.
  • The 1542 Tulane elevator contractor has a crew of 4 men working on the main elevators in the building; the bank located in the rear hall.
    • Repair work also continues on the freight elevator.
  • Conduit has been installed for the electrical wiring to the new disconnects serving the new basement sump pumps.
    • The new disconnects are being installed in Room 117B.
    • The three new submersible sump pumps have been installed.
  • A leaking fan coil in Room A328 has been capped off.
  • Miscellaneous leaks are being repaired as the new pumps are tested and the plumbing system is returned to normal operations.
    • Some drains in the basement have been clogged and are being cleaned out.
  • The water treatment contractor inspected the chemical feed pots for the chilled water and heating water loops.
  • FEMA team members have been on site examining some of the additional damage in the 1542 Tulane basement as part of the scope alignment process.

Project to Make Temporary Repairs to the Dental School:

  • The sub-contractor providing the dehumidification units at the Dental School changed out all of the air monitors on the units to analyze the air in the building. 
    • The units are keeping the humidity at approximately 40% which is sufficient to keep mold from growing.
  • As needed repair work continues on the various air handling units in the Dental School Clinic and Administration Buildings.
  • Repairs are being made to thermostats and zone actuators throughout the Clinic and Administration Buildings.
  • Exposed ducts in the basement and 1st floor are being capped off.
  • The inspection and assessment of the boilers by the manufacturer to determine the extent of the storm damage is underway. The assessment will help to determine the repairs that may be needed to get the boilers back into operation.
  • The elevator contractor continues to make needed repairs to the elevators.
    • Repair personnel worked over the weekend of 12/9 and 12/16 installing parts for the freight elevator. Repairs underway or completed:
      • The exterior elevator doors in the basement have been replaced.
      • The new gates on the elevator itself as well as the counterweights and the rail have been installed.
  • All of the parts necessary for the repair of the flood gate have arrived. Repairs are underway.
  • The shut off valves on the four sump pumps in the basement have been installed.
  • The contractor started running the water in the buildings to clear the system and to check the water pressure.
  • All of the old copper pipe has been removed from the 1st floor.
  • A specialty contractor inspected the tunnel connecting the Power Plant to the Dental School basement to check for the presence of any asbestos and the presence of any chlorine in the piping.

Project to Provide New Lighting for the Student Parking Lots:

  • Parking Lot #2:  The new lights are working, the fence has been restored and the clean up of the site is complete.
  • Parking Lot #3: The installation of the new lighting fixtures and the related electrical connections has been completed.  The clean up of the site is complete.
  • The new lights are all working in Student Lot #2 and #3. The final adjustments of the new light fixtures to reduce shadows and provide even illumination throughout the parking lots have been completed.

Miscellaneous Work:

  • The de-ionized water system in the CSRB has been sanitized.
    • The contractor put a 35% hydrogen peroxide solution into the system, changed the final filters, ran the de-ionized water system in the labs to clean out the delivery system, and performed testing to confirm that the sanitizing was complete.
  • The atrium on the 3rd floor of the MEB is undergoing an exciting renovation. Within a few weeks, we hope to produce a welcoming and comfortable multi-function space where students can gather to study, socialize, and eat. The atrium will be divided into 2 main areas. At the entry will be a spacious lounge area with contemporary sectional sofas, plush club chairs, and occasional tables. Beyond the lounge will be an updated eating area with ample seating at café-style tables. The space will also offer wireless technology, untethering laptops from the perimeter of the room. We expect that the project will take approximately 30 days to complete.
    • The old ceiling tiles and grid have been removed from the atrium. 
    • The old ceiling tiles and grid was completely removed from the hallways. New grid has been installed in the connecting hallways in the MEB.
    • All of the HVAC duct work has been raised.  New connections are being made.
    • All of the grilles and flex connections have been removed.
    • All return grilles and supply grilles were removed from the 3rd floor MEB hallways.
    • All of the concrete floor demolition is complete.
    • The old ceramic wall tile has been removed. 
    • The de-ionized water line to Allied Health has been relocated.

We encourage you to follow the details of our progress through the daily updates which are posted on the web site.